Exercise 6 Solve the IVP y +54 + 6y = u(t - 1) +(t – 2), y(0) = 0, y(0) = 1
Answer #1

Kindly go through the solution and let me know in case of any doubt or further clarification is needed in the comment box.
Thanks for the question :)
And your upvote will be really appreciable ;)
y + 5y +6y - Ult-1) + d(+-2) 1101=0, vlovi L(4)= syis) - Y(0)= SYS) [(y)= $? yls) - SV10) -YO) = 3? yes) -1 syis) - 1 + 55